how long must a guy wire be to run from the top od a 20-ft pole to a point on the ground 8ft from the basse of the pole?
Answer by stanbon(48558)

how long must a guy wire be to run from the top od a 20-ft pole to a point on the ground 8ft from the basse of the pole?
--------------
Draw the picture.
You have a right triangle with a leg of 20 and a leg of 8.
You want to know the hypotenuse.
Use Pythagoras:
h^2 = 20^2 + 8^2
h^2 = 464
h = 21.54... ft
================
Cheers,
Stan H.